CHAPTER 7 K-theory

7.1. Vector bundles 7.2. The ring K(X)

7.3. Chern-Weil theory and the Chern character

7.4. K_{1}(X) and the odd Chern character

7.5. C^{∗} algebras 7.6. K-theory of an algeba

7.7. The norm closure of Ψ^{0}(X)

7.8. The index map 7.9. Problems

126 7. K-THEORY